OpenStudy (anonymous):

Find equation of a line that passes through (-2, 8) and is perpendicular to x+3y-8=0 PLEASE HELP FINAL EXAM TOMORROW

Nnesha (nnesha):

first solve this equation  solve for y

Nnesha (nnesha):

add 8 both side first

OpenStudy (anonymous):

3y=8-x

Nnesha (nnesha):

divide both side by 3

OpenStudy (anonymous):

y=8/3-x/3

Nnesha (nnesha):

yes right so what is slope of your equation

OpenStudy (anonymous):

8/3

Nnesha (nnesha):

slope intercept equation is y = mx + b where m is slope ALWAYS wit x and b is y-intercept

OpenStudy (anonymous):

is 8/3 not the slope?

Nnesha (nnesha):

nope like i said slope is m which is always with x so 8/3 not with x

OpenStudy (anonymous):

1/3?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

-1/3 i mean?